The problem is:
The top of the John Hancock Building in Chicago is a rectagle whose lenth is 60 ft more than the width. The perimeter is 520 ft. Find the width and the length of the rectangle. Find the area of the rectangle.
At first, I though i had the answer, but i didn't look to see that the perimeter was 520 ft. so i just got lost and didn't know where to start anymore. so far, i got: let x =width ; let x=60 = length
x+(x+60) = 520 ft
but when i worked it out, it didn't seem right. please help! =)

Answer by mathrebel(25)

Your reasoning is correct, but your formula is not. The perimeter of a rectangle is is 2*(l+w), so the correct formula is:
2*(x+(x+60))=520
2*(2x+60)=520
divide by 2;
(2x+60)=260
subtract 60,
2x=200
divide by two,
x=100=width
length is 60+x=160
The area is l*w=100*160=16000